OpenStudy (anonymous):

make k the subject of the following two equation, showing each step of your working 1a) 4t=k/7 - 3 1b) 3k+11t= 3+7kt

OpenStudy (anonymous):

b) \[3k-7kt=3-11t \] k(3-7t)=3-11t k=(3-11t)/(3-7t)

OpenStudy (anonymous):

a) 4t = k/7-3 4t = k/4 16t=k

OpenStudy (anonymous):

thanks